Luis and Cameron shared some stickers is the ratio 5:2.
Luis received 15 more stickers than Cameron.
How many stickers were there altogether?

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52782)

.
The condition means that Luis has 5x stickers, while Cameron has 2x stickers, where x is some number, which is 
the common measure of the two amounts.


From the condition,  


    5x - 2x = 15,


which implies


    3x = 15  and  x = 15/3 = 5.


Thus Luis has 5*5 = 25 stickers, while Cameron has 2*5 = 10 stickers.


Answer by greenestamps(13200)



Given the ratio of 5:2, let the two numbers of stickers be 5x and 2x. It is a common way to start any problem involving ratios.

The difference between the two amounts is 5x-2x = 3x; and we know that difference is 15.

3x=15 means x=5.

So the numbers of stickers the two of them have are 5x=25 and 2x=10; the total number of stickers is 25+10 = 35.

Or you could simply say the total number of stickers is 5x+2x = 7x = 35.